OpenStudy (anonymous):

find the derivative of f(x) = 4 / (3x - x^2)

OpenStudy (anonymous):

\[f(x)=\frac{ 4 }{ 3x-x^2 }\]

OpenStudy (anonymous):

\[\frac{ 4 }{ 3(x+h)-(x+h)^2 }-\frac{ 4 }{ 3x-x^2 }\]

OpenStudy (anonymous):

are you using the rules for derivatives or are you still using the limit method?

OpenStudy (anonymous):

limit method

OpenStudy (anonymous):

ahh its a lot more complicated that way unfortunately give me a sec to write out the work myself to make sure I got it then I'll walk through the process with you

OpenStudy (anonymous):

is it just 0, if i use substitution?

OpenStudy (anonymous):

wait no, that's just the limit, i need the derivative

OpenStudy (anonymous):

It's definitely not zero. If you have an answer you want me to check I can do that a lot quicker.

OpenStudy (anonymous):

which limit method are you using there are two.

OpenStudy (anonymous):

umm, i didnt learn the one with deltas, but...\[\lim(x>0) \frac{ f(x+h)-f(x) }{ h }\]

OpenStudy (anonymous):

are you sure that your formula is not suppose to have h approaching to zero not x?

OpenStudy (anonymous):

oh yeah sorry lim(h->0)

OpenStudy (anonymous):

so lets write out that formula subbing in the f(x) part.\[\lim_{h \rightarrow 0}\frac{ \frac{ 4 }{ 3(x+h)-(x+h)^2 }-\frac{ 4 }{ 3x-x^2 } }{ h }\]

OpenStudy (anonymous):

rationalize denominators?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\frac{ 4(3x-x^2)-4[3(x+h)-(x+h)^2]}{ (3x-x^2)[3(x+h)-(x+h)^2] }\]

OpenStudy (anonymous):

the numerator should now be \[\frac{ 4(3x-x^2)-4(3x+3h-x^2-2hx-h^2) }{ (3x-x^2)(3x+3h-x^2-2hx-h^2) }\]

OpenStudy (anonymous):

now expand

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

the numerator I mean we may luck out with the denominator

OpenStudy (anonymous):

when i tried the problem before, i got stuck in the denominator algebra

OpenStudy (anonymous):

numerator: \[-12h + 8hx +4h^2\]

OpenStudy (anonymous):

in fact we do because we have the numerator as \[\frac{ 4(h^2-3h+2hx) }{ (3x-x^2)(3x+3h-x^2-2hx-h^2 }\]

OpenStudy (anonymous):

or should i factor the 4 out?

OpenStudy (anonymous):

It doesn't truly matter. Either way works fine. It's more standard to factor it out if there is a common factor to pull out.

OpenStudy (anonymous):

now for messy denominator

OpenStudy (anonymous):

What is more important is that you recognize we can pull an h out now

OpenStudy (anonymous):

DON"T DO THAT

OpenStudy (anonymous):

IIt's a giant waste of time

OpenStudy (anonymous):

yeah, but I need to distribute 3x-x^2

OpenStudy (anonymous):

no you don't. Wait for me to type out the full equation one more time and you will see what I mean.

OpenStudy (anonymous):

\[\lim_{h \rightarrow 0}\frac{ \frac{ 4h(h-3-2x) }{ (3x-3x^2)(3x+3h-x^2-2hx-h^2 } }{h }\]

OpenStudy (anonymous):

notice the cancellation we can preform now?

OpenStudy (anonymous):

The "h" in the very top of the question cancels with the "h" in the very bottom of the question.

OpenStudy (anonymous):

I think:well i did it anyways\[9x+9xh-3x^3-6hx^2-3xh^2-3x^3-3hx^2+x^4+2hx^3+h^2x^2\]

OpenStudy (anonymous):

so the h on the bottom is really h/1, soo we can multiply the top by 1/h right?

OpenStudy (anonymous):

if thats the way you want to think of it sure! The important part is that they cancel.

OpenStudy (anonymous):

Since they cancel each other we can stop and ask ourselves if we let h approach zero is there still a problem? Do you think there is?

OpenStudy (anonymous):

well, how does the bottom h cancel all the others?

OpenStudy (anonymous):

nvm

OpenStudy (anonymous):

I guess you see now how the rest of the hs do not cause the denominator to approach zero itself. What is the final answer then? Do try to simplify!

OpenStudy (anonymous):

\[\frac{ 2(-3h+2hx+4h^2) }{ 3x-x^2 }\]

OpenStudy (anonymous):

because h approaches 0

OpenStudy (anonymous):

u can "ignore" h multiplied by any term

OpenStudy (anonymous):

\[\frac{ 2 }{ (3x-x^2) }\]

OpenStudy (anonymous):

not quite

OpenStudy (anonymous):

you are making a mistake in your simplification. Instead of ignoring the h as you put it actually sub in zero everywhere you see the letter h and see how that works for simplification.

OpenStudy (anonymous):

\[\frac{ 0 }{ 3x-x^2 }\]

OpenStudy (anonymous):

nope here I'll write in the sub for you

OpenStudy (anonymous):

\[\frac{ 4(0-3+2x) }{(3x-x^2)(3x+0-x^2-0-0) }\]

OpenStudy (anonymous):

I think I switched the signs on the numerator though

OpenStudy (anonymous):

crap we seem to have messed up somewhere one sec

OpenStudy (anonymous):

nope we are fine! I checked our final answer with a math program on my laptop.

OpenStudy (anonymous):

ok thanks!

OpenStudy (anonymous):

\[\frac{ 4(2x-3) }{ (3x-x^2)^2 }\]

OpenStudy (anonymous):

wait isnt the denominator 2(3x-x^2), and if it is, then we can factor out the 4 to a 2